isabel is 13 years older than marie. in 5 years, isabel age will be 10 less than twice mai's age. how old is each now?
Answer by Alan3354(69443)

isabel is 13 years older than marie. in 5 years, isabel age will be 10 less than twice mai's age. how old is each now?
===========
I = M + 13
In 5 years:
I = 2M - 10
-------
Sub for I
M + 13 = 2M - 10
M = 23 (in 5 years)
etc